# Searching for Exotrojans in Pulsar Binary Systems

### v1

@@ Appendix B Upper Limits on the Trojan Mass
To turn an UL on $\mathcal{T}$ into an upper limit on $m_{\rm T}$, some assumptions have to be made about the population of undetected Trojans in the pulsar binary system, as well as the pulsar and its dominant companion. Here, we only concern ourselves with the most massive of such Trojans. We start with the definition of $\mathcal{T}=\alpha z\delta(1-\delta)$. Defining $\epsilon=m_{\rm T}/m_{\rm c}$ and ignoring $\mathcal{O}(\epsilon^{2})$ terms, we find
@@
To find the 95% UL on $m_{\rm T}$, we sample $\mathcal{T}$, $m_{\rm c}$, $K$, and $z$ from their expected distributions in a Monte Carlo fashion. $\mathcal{T}$ is sampled from the reweighted marginalized posterior (see Section 4.2). $K$ is fixed to the value of $a_{\rm p}\sin{I}$ determined in the timing model because its uncertainty is extremely low compared to the other parameters in Equation B3. While pulsar Trojans clearly do not share the same evolutionary history as their Jovian brothers, we choose to sample $z$ from the distribution of Jupiter Trojans with inclinations less than 10° as shown in Figure 3 of F. Marzari et al. (2003). $m_{\rm c}$ is sampled indirectly by solving the binary mass function (Equation 17) for each sample of $m_{\rm p}$ and $\sin{I}$. We choose the standard $\cos{I}\sim\mathcal{U}(0,\,1)$, which assumes a random orientation on the unit sphere of the vector orthogonal to the orbital plane of the pulsar and companion. We choose $m_{\rm p}$ from the bimodal pulsar mass function found by J. Antoniadis et al. (2016).

### v2

@@ Appendix B Upper Limits on the Trojan Mass
To turn an UL on $\mathcal{T}$ into an upper limit on $m_{\rm T}$, some assumptions have to be made about the population of undetected Trojans in the pulsar binary system, as well as the pulsar and its secondary. Here, we only concern ourselves with the most massive of such Trojans. We start with the definition of $\mathcal{T}=\alpha z\delta(1-\delta)$. Defining $\epsilon=m_{\rm T}/m_{\rm c}$ and ignoring $\mathcal{O}(\epsilon^{2})$ terms, we find
@@
To find the 95% UL on $m_{\rm T}$, we sample $\mathcal{T}$, $m_{\rm c}$, $K$, and $z$ from their expected distributions in a Monte Carlo fashion. $\mathcal{T}$ is sampled from the reweighted marginalized posterior (see Section IV.2). $K$ is fixed to the value of $a_{\rm p}\sin{I}$ determined in the timing model because its uncertainty is extremely low compared to the other parameters in Equation B3. While pulsar Trojans clearly do not share the same evolutionary history as their Jovian brothers, we choose to sample $z$ from the distribution of Jupiter Trojans with inclinations less than 10° as shown in Figure 3 of Marzari et al. (2003). $m_{\rm c}$ is sampled indirectly by solving the binary mass function (Equation 28) for each sample of $m_{\rm p}$ and $\sin{I}$. We choose the standard $\cos{I}\sim\mathcal{U}(0,\,1)$, which assumes a random orientation on the unit sphere of the vector orthogonal to the orbital plane of the pulsar and companion. We choose $m_{\rm p}$ from the bimodal pulsar mass function found by Antoniadis et al. (2016).